Suppose we wanted to build a 94% confidence interval to estimate the true mean retirement age for Canadian sailors using a sample size of n = 100. What critical value would be used?
Answer #1

Solution :

Given that ,

n = 100

df = 100 - 1 = 99

/ 2 = 0.06 / 2 = 0.03

t /2,df = 1.903

Critical value = 1.903